The specific heat of aluminum is 0.902j/g.°C.How much heat is absorbed by an aluminum pie tin with a mass of 473g to raise its temperature from room temperature (23.00°C)to oven temperature(375°F


1
Expert's answer
2021-10-07T00:55:01-0400

375oF =(375-32)*(5/9)oC


=190.5oC


Increase in temperature =190.5-23


=167.5


Required Heat= 0.902*167.5*473 J


=71463.205 J